OpenStudy (anonymous):

where did the 4 come from

OpenStudy (bahrom7893):

where?

OpenStudy (bahrom7893):

it fell off the sky... 2+2 is 4; 2*2 is 4

OpenStudy (anonymous):

come from 2+2 = 4 ; 2^2=4 : -(-2-2)=4

OpenStudy (bahrom7893):

it appeared out of nowhere

OpenStudy (bahrom7893):

0-(-4) = 4

OpenStudy (anonymous):

something do with some dude about to be hit on the head with a golf ball

OpenStudy (greencat):

It seems like the 4 came from "4 years ago".